Question 20 Calculate the present value of 20,000 monetary units if they are received in 15 months, assuming an APR of 5% with quarterly compounding (round to the nearest integer).

hello guys may i get a proper explanation for solving this thank you
Answer #1

15 months is 5 quarters (15/3=5)
We know present value of an amount A received in quarter t at rate r compounded quarterly is A/(1+r/4)^t
Hence, here Present Value=20000/(1+5%/4)^5=18795.54124